A student is using the elimination method to solve the system of equations below. What is the best first step?

3x - y = 5
2x + 3y = 10

Answer by addingup(3677)

3x - y = 5
2x + 3y = 10
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3x - y = 5
Multiply times 3:
9x-3y = 15
Now add both equations:
9x-3y = 15
+
2x+3y = 10
-----------
11x = 25
x = 2.27
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3(2.27)-y= 5
6.81-y = 5 Multiply both sides times -1
-6.81+y= -5 Add 6.81 on both sides:
y = 1.81
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
3(2.27)-1.81= 5
6.82-1.81= 5 Subtract on left, you get 5 We have the correct answer.
The other equation:
2x+3y = 10
2(2.27)+3(1.81) = 10
4.55+5.45 = 10 Again we have the right answer.
J